# Discrete quantum groups and their duals

### v1

@@ Abstract
Discrete quantum groups were introduced as duals of compact quantum groups by Podleś and Woronowicz in 1990 [Po-Wo]. Shortly after, they were defined and studied intrinsically in [Ef-Ru] and [VD-discrete]. In 1998, with the introduction of the multiplier Hopf algebras with integrals (also called algebraic quantum groups), the duality between discrete and compact quantum groups became part of the more general duality in the self-dual category of these algebraic quantum groups [VD-alg]. Again a few years later the duality was extended to all locally compact quantum groups (see e.g. [Ku-Va2]).
@@
The operator algebra approach to quantum groups finds it origin in the attempts to generalize Pontryagin’s duality theorem for abelian locally compact groups ([Po]) to the non-abelian ones. The first results in this direction all describe a dual of a (not necessarily abelian) locally compact group and show how to recover the original group from this dual object. Unfortunately, the nice symmetry, present in the abelian case, had to be given up. With the concept of a Kac algebra, developed independently by Kac and Vainerman ([Va-Ka]) and Enock and Schwarz ([En-Sw1, En-Sw2]), the duality was again realized within a sefl-dual category of objects.
However, the work of Drinfel’d ([Dr]) and Jimbo ([Ji]) made people realize that one of the conditions, imposed on the antipode in the theory of Kac algebras, was too restrictive. At the same time, Woronowicz ([Wo1]) constructed the quantum $SU_{q}(2)$ as a compact quantum group within the operator algebra framework. Also this example did not satisfy the Kac algebra axioms. These two events were the starting point for a new search for a self-dual category of objects, strictly larger than the category of Kac algebras so as to include also the compact quantum groups like Woronowicz’ $SU_{q}(2)$.
The result is now known as the theory of locally compact quantum groups. The development of the final object is due to Masuda, Nakagami and Woronowicz on the one hand ([Ma-Na, Ma-Na-Wo]) and Kustermans and Vaes on the other hand ([Ku-Va1, Ku-Va2, Ku-Va3], see also [VD-kerala, VD-sigma].
An intermediate step towards this result came with the theory of multiplier Hopf -algebras with positive integrals (sometimes called -algebraic quantum groups) ([VD-mha, VD-alg]). It is a category of objects, containing both the discrete and the compact quantum groups and it is also self-dual. It is not general enough to contain all locally compact quantum groups. On the other hand however, it is relatively simple and purely algebraic and the theory is rich enough to incorporate most algebraic features of the general theory, avoiding the more difficult technicalities.
The theory of locally compact quantum groups is now well-developed. It is well-understood and there are interesting and plenty of examples, and there are the common general construction methods. And although there is now also a simplified approach to the theory ([VD-sigma]), it remains technically quite involved and relies heavily on Tomita-Takesaki theory techniques. This is probably one of the reasons why it seems that presently, the theory is not very well absorbed by the mathematical community. Perhaps also the fact that there are so many differences in conventions used by various authors (see the appendix).
@@
In the first place, we give a somewhat new and updated approach of the theory of discrete quantum groups. As mentioned already, discrete quantum groups were first introduced by Podleś and Woronowicz as duals of compact quantum groups ([Po-Wo]). Later, discrete quantum groups were intrinsically defined and studied by Effros and Ruan ([Ef-Ru]) and in [VD-discrete]. Therefore, we now look at the work of Podleś and Woronowicz as a way to show that the dual of a compact quantum group (or rather a compact quantum matrix group) is a discrete quantum group. In this sense, their result is a special case of the more general duality between locally compact quantum groups, in particular, between -algebraic quantum groups. Here, undoubtedly, the general case is simpler than the special case as it is treated in the paper by Podleś and Woronowicz. The extra property that the underlying algebra of the dual is a direct sum of matrix algebras follows in the first place from results within the theory of compact quantum groups, but it can also be obtained from the existence of the cointegral in the dual.
@@
In Section 2 we apply the results obtained in the first section to the special case of a discrete quantum group. This is defined as a multiplier Hopf -algebra $(A,\Delta)$ of discrete type with the extra constraint that $A$ is a direct sum of full matrix algebras $A_{\alpha}$ over $\mathbb{C}$, see Definition 2.2. In this case, the left cointegral can be constructed from the axioms of a multiplier Hopf -algebra whose underlying -algebra is such a direct sum. One of the main results here is that the square $S^{2}$ of the antipode $S$ leaves the components invariant. It is a homomorphism and has to be implemented by and invertible element on each component $A_{\alpha}$. As it turns out, it is in fact implemented by a positive invertible element $q_{\alpha}$ in $A_{\alpha}$. The integrals on $A$ are expressed in terms of the standard traces on the components, using these elements $q_{\alpha}$. Hence, we see that the integrals and the corresponding data, are all expressible using the data of the separability idempotent $\Delta(h)$. This treatment of discrete quantum groups is somewhat different from the original way as it is done in [VD-discrete].
@@
In Section 4, we simply discuss the classical cases of a discrete group and of the dual of a compact group. This is just to illustrate the preceding treatment of discrete quantum groups. In a separate paper [VD-suq2] we give a detailed treatment of the discrete quantum group $su_{q}(2)$ as another example of the theory of discrete quantum groups as it is studied in this paper. I believe that this is new although the underlying ideas are not. Indeed, in earlier work the discrete quantum group $su_{q}(2)$ has been studied as the dual of the compact quantum group $SU_{q}(2)$ as introduced by Woronowicz in [Wo1], see again [Po-Wo]. In [VD-suq2] we reverse the order and start with the discrete quantum group. The dual is then the compact quantum group $SU_{q}(2)$.
@@
when $a$ is an element of $A$. This can be done, provided we multiply in all but one factor with elements from $A$, left or right. The use of the Sweedler notation for multiplier Hopf algebras has been introduced in [Dr-VD]. The technique has been refined in [VD-tools] and in an appendix of [VD-lnalg]. For these discrete quantum groups, it is relatively transparent as one can simply multiply with central projections, in other words, one can restrict to components.
@@
The standard references for the theory of Hopf algebras are [Ab] and [Sw]. There is also the more recent work of Radford [Ra1].
The main references for multiplier Hopf algebras and algebraic quantum groups are [VD-mha] and [VD-alg]. Other possible references here are [VD-Zh1] and [VD-part1, VD-part2, VD-part3], as well as [Ti]. See also the more recent paper [DC-VD] with some new results. For discrete quantum groups we refer to [Po-Wo]], [Ef-Ru] and [VD-discrete] but our treatment here will be closest to [VD-discrete]. For compact quantum groups we have the basic references [Wo2] and [Wo3], see also [Ma-VD] for an other treatment.
The main references for the theory of locally compact quantum groups that we use are [Ku-Va2] and [Ku-Va3]. See also the newer and simplified approach in [VD-sigma].
@@ Acknowledgments
It can happen that $\varepsilon(h)=0$ and that $h$ and $k$ are different. Still, also in that case, one can show that cointegrals are unique, see [VD-Zh0].
@@
We will now prove several properties of cointegrals. Most of them can be found already in [VD-Zh0].
@@
This result, as well as some of the following results, are also found in [VD-Zh0].
@@
This is Radford’s formula proven in [Ra0] for the fourth power of the antipode because $\widehat{\delta}=1$. That $\widehat{\delta}=1$ follows from the general theory. Indeed, we know that $\langle a,\widehat{\delta}\rangle=\varepsilon(\sigma^{-1}(a))$, see e.g. Proposition 2.3 in [VD-part1] or Proposition 1.2.3 in [VD-lnalg]. In this case we have $S^{2}=\sigma$ and because $\varepsilon(S^{-2}(a))=\varepsilon(a)$ we get $\langle a,\widehat{\delta}\rangle=\varepsilon(a)$ and therefore $\widehat{\delta}=1$.
@@
  • Proposition The element $\Delta(h)$ is a (regular) separability idempotent in $M(A\otimes A)$ in the sense of [VD-si1].
  • Proof: i) First we look at the setting in [VD-si1]. For the algebras $B$ and $C$ we take $A$ and we let $E=\Delta(h)$. The algebra $A$ is non-degenerate and $\Delta(h)$ is an idempotent in $M(A\otimes A)$. Because $A$ is a regular multiplier Hopf algebra we know that elements of the form $\Delta(h)(1\otimes a)$ and $\Delta(h)(a\otimes 1)$ belong to $A\otimes A$ for all $a\in A$ and similarly with $a$ on the other side. This shows that Assumption 1.1 of [VD-si1] is fulfilled and that $\Delta(h)$ is regular in the sense of that assumption.
    ii) We have shown that the legs of $\Delta(h)$ are all of $A$ and hence it is full. Also Proposition 1.4 in [VD-si1] here says that $\Delta(h)(1\otimes a)=0$ only if $a=0$, and similarly $(a\otimes 1)\Delta(h)=0$ only if $a=0$. This we knew already because we have a multiplier Hopf algebra and $h\neq 0$.
    iii) Finally we come to Definition 1.5 of [VD-si1]. Since we know here already that
@@
The maps $S:B\to C$ and $S^{\prime}:C\to B$ we have in Proposition 1.7 of [VD-si1] are precisely the antipode $S$ of $A$. If we look at the Proposition 1.8 in [VD-si1], we get here
@@
  • Remark In the case where the left cointegral $h$ satisfies $\varepsilon(h)=1$, so that $\Delta(h)$ is a separability idempotent, we can apply general results from [VD-si1] to obtain the existence of $\varphi$ and $\psi$, as in Theorem 1.11 above. However, we only would get $(\iota\otimes\varphi)\Delta(h)=1$ and $(\psi\otimes\iota)\Delta(h)=1$ and we would still need the argument given in the proof of Theorem 1.11 to obtain left and right invariance on all of $A$.
    Also the results in Proposition 1.13 and Proposition 1.16 already follow from the general theory in [VD-si1].
In the next section, we will give an explicit formula for $\Delta(h)$ and refer to this item on separability idempontents.
@@
The above definition of a coproduct is also possible when $A$ is any -algebra with a non-degenerate product (see [VD-mha]). If we do not assume the conditions in i), but only assume that $\Delta$ is non-degenerate (i.e. that $\Delta(A)(A\otimes A)=A\otimes A$), it is still possible to give a meaning to coassociativity, see e.g. the note on coassociativity [VD-coass, VD-coass1]. This is not so hard to see in the case of a direct sum. However, as we will see, the conditions in i) are quite natural and similar conditions are often taken as part of the notion of a coproduct itself.
@@
In Section 4 we also consider some other important special cases and we refer to [VD-suq2] for a more specific example.
@@
    So, the pair $(A,\Delta)$ is a multiplier Hopf -algebra as defined in [VD-mha].
    ii) Conversely, if $(A,\Delta)$ is a multiplier Hopf -algebra and if $A$ is a direct sum of matrix algebras, then it is a discrete quantum group in the sense of Definition 2.2. The reason is that in any multiplier Hopf -algebra, there exists a counit and an antipode (see [VD-mha].
So, a discrete quantum group is nothing else but a multiplier Hopf -algebra where the underlying algebra is a direct sum of matrix algebras (cf. the definition given of a discrete quantum group in [VD-discrete]).
@@
  • Proposition When $(A,\Delta)$ is a discrete quantum group, then it is a multiplier Hopf algebra of discrete type. The left contegral $h$ satisfies $\varepsilon(h)=1$, it is an idempotent and $S(h)=h$. It is also a right cointegral.
@@
The maps $a\mapsto S(a)^{*}$ and $a\mapsto S^{2}(a)$
There are plenty of traces on the algebra $A$, but on each component, the trace is unique up to a scalar. We will further work with the standard traces. Recall that $A_{\alpha}$ is the algebra of $n(\alpha)\times n(\alpha)$ matrices over the complex numbers.
@@
  • Notation Denote by $\omega_{\alpha}$ the unique trace on $A_{\alpha}$ satisfying $\omega_{\alpha}(1_{\alpha})=n(\alpha)$. We use $\omega$ for the trace on $A$ with these components on $A_{\alpha}$.
2.10.
  • Notation Denote $q=(\omega\otimes\iota)\Delta(h)$ and $q_{\alpha}$ for the components of $q$. Because of Proposition 2.8 we have $q_{\alpha}=(\omega_{\overline{\alpha}}\otimes 1)\Delta(h)$.
The element $q$ is defined in the multiplier algebra $M(A)$.
2.11.
  • Proposition The element $q$ is positive and invertible in $M(A)$. Further $aq=qS^{2}(a)$ for all $a\in A$.
@@
    ii) For $a\in A_{\alpha}$ we get $aq_{\alpha}=q_{\alpha}S^{2}(a)$. Because $q_{\alpha}\neq 0$ we have that it is invertible. Because $\Delta(h)$ is positive and the trace is positive, we have that $q$ is positive.
In a similar way we can consider $(\iota\otimes\omega)\Delta(h)$. It will be a positive invertible element $p\in M(A)$ satisfying $ap=S^{-2}(a)p$ for all $a\in a$.
It is natural for $S^{2}$ to be implemented by a self-adjoint invertible element because of the fact that $S^{2}(a)^{*}=S^{-2}(a^{*})$ for all $a$, but that it can be implemented by a positive invertible elements is not automatic.
We will see later that $\omega_{\alpha}(q_{\alpha})=1$.
The scaling group and the unitary antipode
It has been shown in [Ku] that algebraic quantum groups admit an analytic structure. In particular, there is a one-parameter group $(\tau_{t})$ of automorphisms of $A$ with the property that all elements in $A$ are analytic. It is characterized by the fact that $S^{2}=\tau_{-i}$. This is called the scaling group. There is also a unitary antipode $R$ such that $S=R\tau_{-\frac{i}{2}}$. This is thought of as the polar decomposition of the antipode. These results are also found in [Ku-VD] and in the more recent survey paper [VD-part3]. The polar decomposition of the antipode is also a feature of the more general locally compact groups, see the references given in the corresponding item in the introduction.
@@
  • Proposition We get $\tau_{t}(a)=q^{-it}aq^{it}$ for all $a\in A$.
  • Proof: From $S=R\tau_{-\frac{i}{2}}$ we get $S^{2}=\tau_{-i}$. We use that $R^{2}=\iota$ and that $R$ and $\tau_{t}$ commute. Then, because $S^{2}(a)=q^{-1}aq$, we must have $\tau_{t}(a)=q^{-it}aq^{it}$.
The scaling group leaves all components of $A$ globally invariant and $\tau_{t}(a)=q_{\alpha}^{-it}aq_{\alpha}^{it}$ for all $a\in A_{\alpha}$.
The situation is different for the unitary antipode $R$. For this we get the following results. Recall that for each index $\alpha$ there is an index $\overline{\alpha}$ such that $S:A_{\alpha}\to A_{\overline{\alpha}}$.
2.13.
@@
2.14.
  • Proposition There exists a vector $\xi_{0}$ in $\mathcal{H}_{\alpha}\otimes\mathcal{H}_{\overline{\alpha}}$ such that for the corresponding operator $x$ we have $x^{*}x=q_{\alpha}^{-1}$.
  • Proof: We have seen before that $aq_{\alpha}=q_{\alpha}S^{2}(a)$. It follows that $x^{*}x$ must be a scalar multiple of $q_{\alpha}^{-1}$. Because $\Delta(h)$ is non-zero, we have elements $\xi_{0}$ such that $\Delta(h)\xi_{0}\neq 0$ and $x\neq 0$. Therefore there exists a vector $\xi_{0}$ such that actually $x^{*}x=q_{\alpha}^{-1}$.
Consider the polar decomposition of $x$. It is of the form $x=J_{\alpha}q_{\alpha}^{-\frac{1}{2}}$ where $J_{\alpha}$ is a conjugate linear isomorphism of $\mathcal{H}_{\alpha}$ to $\mathcal{H}_{\overline{\alpha}}$.
For the unitary antipode $R$ we have the following.
2.15.
@@
In this item we fix an index $\alpha$ in $I$.
@@
    $\Delta(h)(1_{\overline{\alpha}}\otimes 1_{\alpha})=\sum_{ij}\lambda_{i}\,S(e_{ij})\otimes e_{ji}.$
@@
    We have used that $qa=S^{2}(a)q$ and that $q=\sum_{i}\lambda_{i}e_{ii}$. We see that indeed
@@
    iii) It follows that $H=c\Delta(h)(1_{\overline{\alpha}}\otimes 1_{\alpha})$. It is only natural that this constant is present because we have only used that $q$ implements $S^{2}$ and that is then true for any scalar multiple of $q$. However, we have a specific element $q$ in $A_{\alpha}$, defined as $(\omega\otimes\iota)\Delta(h)$. If we apply the standard trace $\omega$ on the first factor on the left we have
    $(\omega\otimes 1)H=\sum_{ij}\lambda_{i}\,\omega(S(e_{ij}))\otimes e_{ji}=\sum_{i}\lambda_{i}e_{ii}=q.$
    On the right we find $cq$ so that $c=1$ and hence $H=\Delta(h)(1_{\overline{\alpha}}\otimes 1_{\alpha})$. This completes the proof.
The number $c$ in the proof above is nothing else but $\omega_{\alpha}(q)$ and so this is $1$ as we announced already. It gives in fact that every component of $\Delta(h)$ in $A_{\overline{\alpha}}\otimes A_{\alpha}$ is a one-dimensional projection. This result is obtained in the original paper, see Proposition 4.4 in [VD-discrete].
We are now ready for the main property here. We refer to our papers on separability idempotents [VD-si1, VD-si2].

### v2

@@ Abstract
Discrete quantum groups were introduced as duals of compact quantum groups by Podleś and Woronowicz in 1990 [18]. Shortly after, they were defined and studied intrinsically in [6] and [28]. In 1998, with the introduction of the multiplier Hopf algebras with integrals (also called algebraic quantum groups), the duality between discrete and compact quantum groups became part of the more general duality in the self-dual category of these algebraic quantum groups [29]. Again a few years later the duality was extended to all locally compact quantum groups (see e.g. [11]).
@@
The operator algebra approach to quantum groups finds it origin in the attempts to generalize Pontryagin’s duality theorem for abelian locally compact groups ([19]) to the non-abelian ones. The first results in this direction all describe a dual of a (not necessarily abelian) locally compact group and show how to recover the original group from this dual object. Unfortunately, the nice symmetry, present in the abelian case, had to be given up. With the concept of a Kac algebra, developed independently by Kac and Vainerman ([24]) and Enock and Schwarz ([7, 8]), the duality was again realized within a self-dual category of objects.
However, the work of Drinfel’d ([4]) and Jimbo ([9]) made people realize that one of the conditions, imposed on the antipode in the theory of Kac algebras, was too restrictive. At the same time, Woronowicz ([43]) constructed the quantum $SU_{q}(2)$ as a compact quantum group within the operator algebra framework. Also this example did not satisfy the Kac algebra axioms. These two events were the starting point for a new search for a self-dual category of objects, strictly larger than the category of Kac algebras so as to include also the compact quantum groups like Woronowicz’ $SU_{q}(2)$.
The result is now known as the theory of locally compact quantum groups. The development of the final object is due to Masuda, Nakagami and Woronowicz on the one hand ([17, 16]) and Kustermans and Vaes on the other hand ([10, 11, 12], see also [30, 34].
An intermediate step towards this result came with the theory of multiplier Hopf -algebras with positive integrals (sometimes called -algebraic quantum groups) ([27, 29]). It is a category of objects, containing both the discrete and the compact quantum groups and it is also self-dual. It is not general enough to contain all locally compact quantum groups. On the other hand however, it is relatively simple and purely algebraic and the theory is rich enough to incorporate most algebraic features of the general theory, avoiding the more difficult technicalities.
The theory of locally compact quantum groups is now well-developed. It is well-understood and there are interesting and plenty of examples, and there are the common general construction methods. And although there is now also a simplified approach to the theory ([34]), it remains technically quite involved and relies heavily on Tomita-Takesaki theory techniques. This is probably one of the reasons why it seems that presently, the theory is not very well absorbed by the mathematical community. Perhaps also the fact that there are so many differences in conventions used by various authors (see the appendix).
@@
In the first place, we give a somewhat new and updated approach of the theory of discrete quantum groups. As mentioned already, discrete quantum groups were first introduced by Podleś and Woronowicz as duals of compact quantum groups ([18]). Later, discrete quantum groups were intrinsically defined and studied by Effros and Ruan ([6]) and in [28]. Therefore, we now look at the work of Podleś and Woronowicz as a way to show that the dual of a compact quantum group (or rather a compact quantum matrix group) is a discrete quantum group. In this sense, their result is a special case of the more general duality between locally compact quantum groups, in particular, between -algebraic quantum groups. Here, undoubtedly, the general case is simpler than the special case as it is treated in the paper by Podleś and Woronowicz. The extra property that the underlying algebra of the dual is a direct sum of matrix algebras follows in the first place from results within the theory of compact quantum groups, but it can also be obtained from the existence of the cointegral in the dual.
@@
In Section 2 we apply the results obtained in the first section to the special case of a discrete quantum group. This is defined as a multiplier Hopf -algebra $(A,\Delta)$ of discrete type with the extra constraint that $A$ is a direct sum of full matrix algebras $A_{\alpha}$ over $\mathbb{C}$, see Definition 2.2. In this case, the left cointegral can be constructed from the axioms of a multiplier Hopf -algebra whose underlying -algebra is such a direct sum. One of the main results here is that the square $S^{2}$ of the antipode $S$ leaves the components invariant. It is a homomorphism and has to be implemented by and invertible element on each component $A_{\alpha}$. As it turns out, it is in fact implemented by a positive invertible element $q_{\alpha}$ in $A_{\alpha}$. The integrals on $A$ are expressed in terms of the standard traces on the components, using these elements $q_{\alpha}$. Hence, we see that the integrals and the corresponding data, are all expressible using the data of the separability idempotent $\Delta(h)$. This treatment of discrete quantum groups is somewhat different from the original way as it is done in [28].
@@
In Section 4, we simply discuss the classical cases of a discrete group and of the dual of a compact group. This is just to illustrate the preceding treatment of discrete quantum groups. In a separate paper [41] we give a detailed treatment of the discrete quantum group $su_{q}(2)$ as another example of the theory of discrete quantum groups as it is studied in this paper. I believe that this is new although the underlying ideas are not. Indeed, in earlier work the discrete quantum group $su_{q}(2)$ has been studied as the dual of the compact quantum group $SU_{q}(2)$ as introduced by Woronowicz in [43], see again [18]. In [41] we reverse the order and start with the discrete quantum group. The dual is then the compact quantum group $SU_{q}(2)$.
@@
when $a$ is an element of $A$. This can be done, provided we multiply in all but one factor with elements from $A$, left or right. The use of the Sweedler notation for multiplier Hopf algebras has been introduced in [3]. The technique has been refined in [31] and in an appendix of [40]. For these discrete quantum groups, it is relatively transparent as one can simply multiply with central projections, in other words, one can restrict to components.
@@
The standard references for the theory of Hopf algebras are [5] and [22]. There is also the more recent work of Radford [20].
The main references for multiplier Hopf algebras and algebraic quantum groups are [27] and [29]. Other possible references here are [26] and [35, 36, 37], as well as [23]. See also the more recent paper [1] with some new results. For discrete quantum groups we refer to [18]], [6] and [28] but our treatment here will be closest to [28]. For compact quantum groups we have the basic references [42] and [44], see also [15] for an other treatment.
The main references for the theory of locally compact quantum groups that we use are [11] and [12]. See also the newer and simplified approach in [34].
@@ Acknowledgments
It can happen that $\varepsilon(h)=0$ and that $h$ and $k$ are different. Still, also in that case, one can show that cointegrals are unique, see [25].
@@
We will now prove several properties of cointegrals. Most of them can be found already in [25].
@@
This result, as well as some of the following results, are also found in [25].
@@
This is Radford’s formula proven in [21] for the fourth power of the antipode because $\widehat{\delta}=1$. That $\widehat{\delta}=1$ follows from the general theory. Indeed, we know that $\langle a,\widehat{\delta}\rangle=\varepsilon(\sigma^{-1}(a))$, see e.g. Proposition 2.3 in [35] or Proposition 1.2.3 in [40]. In this case we have $S^{2}=\sigma$ and because $\varepsilon(S^{-2}(a))=\varepsilon(a)$ we get $\langle a,\widehat{\delta}\rangle=\varepsilon(a)$ and therefore $\widehat{\delta}=1$.
@@
  • Proposition The element $\Delta(h)$ is a (regular) separability idempotent in $M(A\otimes A)$ in the sense of [32].
  • Proof: i) First we look at the setting in [32]. For the algebras $B$ and $C$ we take $A$ and we let $E=\Delta(h)$. The algebra $A$ is non-degenerate and $\Delta(h)$ is an idempotent in $M(A\otimes A)$. Because $A$ is a regular multiplier Hopf algebra we know that elements of the form $\Delta(h)(1\otimes a)$ and $\Delta(h)(a\otimes 1)$ belong to $A\otimes A$ for all $a\in A$ and similarly with $a$ on the other side. This shows that Assumption 1.1 of [32] is fulfilled and that $\Delta(h)$ is regular in the sense of that assumption.
    ii) We have shown that the legs of $\Delta(h)$ are all of $A$ and hence it is full. Also Proposition 1.4 in [32] here says that $\Delta(h)(1\otimes a)=0$ only if $a=0$, and similarly $(a\otimes 1)\Delta(h)=0$ only if $a=0$. This we knew already because we have a multiplier Hopf algebra and $h\neq 0$.
    iii) Finally we come to Definition 1.5 of [32]. Since we know here already that
@@
The maps $S:B\to C$ and $S^{\prime}:C\to B$ we have in Proposition 1.7 of [32] are precisely the antipode $S$ of $A$. If we look at the Proposition 1.8 in [32], we get here
@@
  • Remark In the case where the left cointegral $h$ satisfies $\varepsilon(h)=1$, so that $\Delta(h)$ is a separability idempotent, we can apply general results from [32] to obtain the existence of $\varphi$ and $\psi$, as in Theorem 1.11 above. However, we only would get $(\iota\otimes\varphi)\Delta(h)=1$ and $(\psi\otimes\iota)\Delta(h)=1$ and we would still need the argument given in the proof of Theorem 1.11 to obtain left and right invariance on all of $A$.
    Also the results in Proposition 1.13 and Proposition 1.16 already follow from the general theory in [32].
In the next section, we will give an explicit formula for $\Delta(h)$ and refer to this item on separability idempotents.
@@
The above definition of a coproduct is also possible when $A$ is any -algebra with a non-degenerate product (see [27]). If we do not assume the conditions in i), but only assume that $\Delta$ is non-degenerate (i.e. that $\Delta(A)(A\otimes A)=A\otimes A$), it is still possible to give a meaning to coassociativity, see e.g. the note on coassociativity [38, 39]. This is not so hard to see in the case of a direct sum. However, as we will see, the conditions in i) are quite natural and similar conditions are often taken as part of the notion of a coproduct itself.
@@
In Section 4 we also consider some other important special cases and we refer to [41] for a more specific example.
@@
    So, the pair $(A,\Delta)$ is a multiplier Hopf -algebra as defined in [27].
    ii) Conversely, if $(A,\Delta)$ is a multiplier Hopf -algebra and if $A$ is a direct sum of matrix algebras, then it is a discrete quantum group in the sense of Definition 2.2. The reason is that in any multiplier Hopf -algebra, there exists a counit and an antipode (see [27].
So, a discrete quantum group is nothing else but a multiplier Hopf -algebra where the underlying algebra is a direct sum of matrix algebras (cf. the definition given of a discrete quantum group in [28]).
@@
  • Proposition When $(A,\Delta)$ is a discrete quantum group, then it is a multiplier Hopf algebra of discrete type. The left cointegral $h$ satisfies $\varepsilon(h)=1$, it is an idempotent and $S(h)=h$. It is also a right cointegral.
@@
The map $a\mapsto S^{2}(a)$ and the scaling group
We consider the implementation of the square $S^{2}$ of the antipode $S$ in the following proposition.
@@
  • Proposition Let $\omega$ be any trace on $A$. Define $q=(\omega\otimes\iota)\Delta(h)$. The element $q$ belongs to $M(A)$ and it satisfies $aq=qS^{2}(a)$ for all $a\in A$. If $\omega$ is positive, then $q$ is positive and if $\omega$ is faithful, then $q$ is invertible in $M(A)$.
@@
    ii) As before, we denote by $q_{\alpha}$ the $\alpha$-component of $q$. When $\omega$ is positive, this will be a positive element in $A_{\alpha}$. When $\omega$ is faithful, all these components will be non-zero. As they implement $S^{2}$, they are invertible. Then also $q$ has an inverse in $M(A)$.
2.10.
  • Remark i) There are plenty of traces on $A$, but on each component, the trace is unique up to a scalar. This implies that the components $q_{\alpha}$ of the element $q$ above are all unique up to a scalar. That is also a consequence of the fact that these components implement the isomorphisms $S^{2}$, restricted to the components.
    ii) Any isomorphism of the algebra of $n\times n$-matrix algebra is implemented by an invertible matrix. Because here $S^{2}(a)^{*}=S^{-2}(a^{*})$ it is easy to see that it can be implemented by a self-adjoint matrix. That actually we get an implementation by a positive matrix is not automatic but appears to be true in this situation. The reason is that already $S$ satisfies $S(a)^{*}=S^{-1}(a^{*})$.
In a similar way we can consider $p=(\iota\otimes\omega)\Delta(h)$ for any trace. Now we have $ap=pS^{-2}(a)$ for all $a\in A$. Again it will be a positive invertible element $p\in M(A)$ if the trace is positive and faithful. Because also $aq^{-1}=q^{-1}S^{-2}(a)$ (where $q$ is the element obtained in the previous proposition), we must have that each component $p_{\alpha}$ of $p$ is a scalar multiple of the corresponding component $q_{\alpha}$ of $q$.
We will now look for the scaling group. Recall that algebraic quantum groups admit an analytic structure, see [14]. In particular, there is a one-parameter group $(\tau_{t})$ of automorphisms of $A$ with the property that all elements in $A$ are analytic. It is characterized by the fact that $S^{2}=\tau_{-i}$. This is called the scaling group.
@@
2.11.
  • Proposition Let $q$ be a positive invertible element in $M(A)$ satisfying $aq=qS^{2}(a)$ for all $a\in A$. Then $\tau_{t}(a)=q^{-it}aq^{it}$ for all $t$ and $a\in A$.
  • Proof: Because $S^{2}(a)=q^{-1}aq$ and $\tau_{-i}=S^{2}$ we must have $\tau_{t}(a)=q^{-it}aq^{it}$.
The scaling group leaves all components of $A$ globally invariant and $\tau_{t}(a)=q_{\alpha}^{-it}aq_{\alpha}^{it}$ for all $a\in A_{\alpha}$. Because $q_{\alpha}$ is unique up to a scalar, the expression on the right does not depend on the choice of $q$ (as expected).
The unitary antipode $R$
The unitary antipode $R$ is characterized so that $S=R\tau_{-\frac{i}{2}}$. This formula is thought of as the polar decomposition of the antipode. These results are also found in [13] and in the more recent survey paper [37]. The polar decomposition of the antipode is also a feature of the more general locally compact groups, see the references given in the corresponding item in the introduction.
For the unitary antipode of a discrete quantum group, we get the following results. Recall that for each index $\alpha$ there is an index $\overline{\alpha}$ such that $S:A_{\alpha}\to A_{\overline{\alpha}}$.
@@
If $q$ is a positive invertible element satisfying $aq=qS^{2}(a)$ we will have that $x^{*}x$ is a scalar multiple of $q_{\alpha}^{-1}$. In particular, given such an element $q$ we can find a vector $\xi$ such that the corresponding $x$ satisfies $x^{*}x=q_{\alpha}^{-1}$.
Then consider the polar decomposition of $x$. It is of the form $x=J_{\alpha}q_{\alpha}^{-\frac{1}{2}}$ where $J_{\alpha}$ is a conjugate linear isomorphism of $\mathcal{H}_{\alpha}$ to $\mathcal{H}_{\overline{\alpha}}$.
2.13.
@@
We consider an element $q$, positive and invertible in $M(A)$ satisfying $aq=qS^{2}(a)$ for all $a\in A$ as before. In this item we fix an index $\alpha$ in $I$ and the component $q_{\alpha}$. It is a positive invertible matrix in $A_{\alpha}$ satisfying $aq_{\alpha}=q_{\alpha}S^{2}(a)$ for all $a\in A_{\alpha}$.
@@
    $\Delta(h)(1_{\overline{\alpha}}\otimes 1_{\alpha})=\frac{1}{c}\sum_{ij}\lambda_{i}\,S(e_{ij})\otimes e_{ji}$
    where $c=\sum_{i}\lambda_{i}$.
@@
    We have used that $aq_{\alpha}=q_{\alpha}S^{2}(a)$ and that $q_{\alpha}=\sum_{i}\lambda_{i}e_{ii}$. We see that indeed
@@
    iii) It follows that $H=c\Delta(h)(1_{\overline{\alpha}}\otimes 1_{\alpha})$.
It is natural that this constant is present because we have only used that $q_{\alpha}$ implements $S^{2}$ and that is then true for any scalar multiple of $q_{\alpha}$. Further observe that indeed, the right hand side does not depend on the choice of $q_{\alpha}$ as it should.
2.15.
  • Notation In what follows we will use $\omega_{\alpha}$ for the standard trace on $A_{\alpha}$, defined by $\omega_{\alpha}(1_{\alpha})=n_{\alpha}$ when $n_{\alpha}$ is the dimension of the matrix algebra $A_{\alpha}$.
Because $S$ is an anti-isomorphism from $A_{\alpha}$ to $A_{(}\overline{\alpha})$ we will have that $\omega_{\overline{\alpha}}\circ S=\omega_{\alpha}$.
We get the following consequence of Proposition 2.14.
2.16.
  • Proposition Each component of $\Delta(h)$ is one-dimensional.
  • Proof: Then
    $(\omega_{\overline{\alpha}}\otimes\omega_{\alpha})\Delta(h)$ $=\frac{1}{c}\sum_{i,j}\lambda_{i}\omega_{\overline{\alpha}}(S(e_{ij}))\omega_{\alpha}(e_{ij})$
    $=\frac{1}{c}\sum_{i}\lambda_{i}=1.$
Observe that this is a result already obtained in the original paper on discrete quantum groups, see Proposition 4.4 in [28].
We are now ready for the main property here. We refer to our papers on separability idempotents [32, 33].